z,w be two complex no.
|z|=|w|
argz+argw=pi
then z equal?

Grade:9

1 Answers

Arun Kumar IIT Delhi
8 years ago
Hello Student,

\\let \\\omega=re^{i\theta} \\=>z=re^{\pi-i\theta}=re^\pi*e^{-i\theta} \\=>z=-\bar\omega

Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty